OpenStudy (anonymous):

What is the sum of a 30-term arithmetic sequence where the first term is 71 and the last term is -103? (3 points) -480 -448 -416 -384

OpenStudy (jumperman):

I believe it's -416, you can check my work, I might be wrong.

OpenStudy (anonymous):

Can u explain please ?

OpenStudy (anonymous):

@phi @SolomonZelman @Solver @mathstudent55

OpenStudy (mathstudent55):

The sum of n terms of an arithmetic series is: \(S_n = \dfrac{n(a_1 + a_n)}2\) where \(S_n\) = sum of n terms \(a_1\) is the first term, and \(a_n\) is the \(n\)th term

OpenStudy (mathstudent55):

\(S_{30} = \dfrac{30[73 + (-103)]}2 \)

OpenStudy (anonymous):

\[S _{30}=-450\]

OpenStudy (anonymous):

Can someone help me finish it please

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Please help me finish this

ganeshie8 (ganeshie8):

first term = 71 , so evaluate below : \(\large S_{30} = \dfrac{30[\color{Red}{71} + (-103)]}2\)

OpenStudy (anonymous):

-480

OpenStudy (anonymous):

Thank You so much @ganeshie8

ganeshie8 (ganeshie8):

yw

OpenStudy (mathstudent55):

Sorry, I wrote 73 for the first term by mistake. It was 71. Answer is -480.